OpenStudy (anonymous):

Simplify square root of 5 open parentheses 10 minus 4 square root of 2 close parentheses. 14 15 square root of 2 5 square root of 2 end root minus 4 square root of 10 None of the above

OpenStudy (calculusxy):

\(\sqrt{5}(10 - 4 \sqrt{2})\) \(10\sqrt{5} - 4\sqrt{5 \times 2}\) \(10\sqrt{5} - 4\sqrt{10}\) this cannot be simplified anymore.

OpenStudy (anonymous):

\[10 \sqrt{5}-4 \sqrt{10} \]

OpenStudy (anonymous):

@robtobey Can you help with this please? http://openstudy.com/study#/updates/5688599ae4b032ed60df42b6 Thanks!